# A Bias–Variance Decomposition for Ensembles over Multiple Synthetic Datasets

### v2

@@ Abstract
Recent studies have highlighted the benefits of generating multiple synthetic datasets for supervised learning, from increased accuracy to more effective model selection and uncertainty estimation. These benefits have clear empirical support, but the theoretical understanding of them is currently very light. We seek to increase the theoretical understanding by deriving bias-variance decompositions for several settings of using multiple synthetic datasets, including differentially private synthetic data. Our theory predicts multiple synthetic datasets to be especially beneficial for high-variance downstream predictors, and yields a simple rule of thumb to select the appropriate number of synthetic datasets in the case of mean-squared error and Brier score. We investigate how our theory works in practice by evaluating the performance of an ensemble over many synthetic datasets for several real datasets and downstream predictors. The results follow our theory, showing that our insights are practically relevant.
@@ 1 Introduction
In supervised learning, van Breugel et al. (2023a) proposed using multiple synthetic datasets in the form of an ensemble. They propose training a predictive model separately on each synthetic dataset, and ensembling these models by averaging their predictions, which they call a deep generative ensemble (DGE). We drop the word “deep” in this paper and use the term generative ensemble, as we do not require the generator to be deep in any sense. The DGE was empirically demonstrated to be beneficial in several ways by van Breugel et al. (2023a), including predictive accuracy, model evaluation, model selection, and uncertainty estimation. Followup work has applied DGEs to improve model evaluation under distribution shifts and for small subgroups (van Breugel et al., 2023b).
@@
$\underbrace{\operatorname*{\mathbb{E}}_{D,y}[(y-g(x;D))^{2}]}_{\mathrm{MSE}}=\underbrace{(f(x)-\operatorname*{\mathbb{E}}_{D}[g(x;D)])^{2}}_{\mathrm{Bias}}+\underbrace{\operatorname*{Var}_{D}[g(x;D)]}_{\mathrm{Variance}}+\underbrace{\operatorname*{Var}_{y}[y]}_{\mathrm{Noise}}$ (1)
where we $f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor. $x$ is considered fixed, so all the random quantities in the decomposition are implicitly conditioned on $x$. While MSE is typically only used with regression problems, the decomposition also applies to the Brier score (Brier, 1950) in classification, which is simply the MSE of class probabilities.
@@
    We derive a bias-variance decomposition for the MSE or Brier score of generative ensembles under an i.i.d. assumption in Theorem 2.2. This decomposition is simply a sum of interpretable terms, which makes it possible to predict how various choices affect the error. In particular, the number of synthetic datasets only affects the variance-related terms in the decomposition, predicting that generative ensembles work best with high-variance and low bias models, like interpolating decision trees.
@@
    We derive several practical considerations from our decomposition, including a simple rule of thumb to select the number of synthetic datasets in Section 2.3. In summary, 2 synthetic datasets give 50% of the potential benefit from multiple synthetic datasets, 10 give 90% of the benefit and 100 give 99% of the benefit. This also applies to bagging ensembles like random forests, which is likely to be of independent interest.
@@
    We evaluate the performance of a generative ensemble on several datasets, downstream prediction algorithms, and error metrics in Section 3. The results show that out theory applies in practice: multiple synthetic datasets generally decrease all of the error metrics, especially for high-variance models where the theory predicts multiple synthetic datasets to have the highest impact. In addition, our rule of thumb makes accurate predictions when the error can be accurately estimated.
@@ 1.1 Related Work
In a similar manner, multiple imputation (Rubin, 1987; Murray, 2018) imputes missing values multiple times to a dataset, which allows correcting uncertainty estimates for the extra uncertainty from missing values. Multiple imputation is typically used with statistical inference in mind, but the imputed datasets could be used to train an ensemble.
@@
Ueda & Nakano (1996) were the first to study the MSE bias-variance decomposition with ensembles, and Gupta et al. (2022); Wood et al. (2023) later extended the ensemble decomposition to other losses. All of these also apply to generative ensembles, but they only provide limited insight for them, as they do not separate the synthetic data generation-related terms from the downstream-related terms.
@@ 2 Bias-Variance Decompositions for Generative Ensembles
In this section, we make our main theoretical contributions. We start by defining our setting in Section 2.1. Then we derive a bias-variance decomposition for generative ensembles in Section 2.2 and derive a simple rule of thumb that can be used to select the number of synthetic datasets in Section 2.3. The generative ensemble decomposition does not apply to several differentially private synthetic data generation methods that generate all synthetic datasets conditional on a single privatised value, so we generalise the decomposition to apply to those in Section 2.4. We also present a third decomposition that applies to Bregman divergences, but is not as informative as the others, in Appendix C.
@@ 2.1 Problem Setting
We assume that the synthetic datasets are generated by first i.i.d. sampling parameters $\theta_{1:m}|D_{r}$ given the real data, and then sampling synthetic datasets $D_{s}^{i}\sim p(D_{s}|\theta_{i})$ given the parameters. This is how synthetic datasets are sampled in DGE (van Breugel et al., 2023a), where $\theta_{i}$ are the parameters of the generative neural network from independent training runs. This also encompasses bootstrapping, where $\theta_{i}$ would be the real dataset [^2] The $\theta_{i}$ random variables are i.i.d. if they are deterministically equal., and $p(D_{s}|\theta_{i})$ is the bootstrapping.
We will also consider a more general setting that applies to some differentially private synthetic data generators that do not fit into this setting in Section 2.4. In Appendix B, we consider settings without the i.i.d. assumption.
@@ 2.2 Mean-squared Error Decomposition
$\begin{split}\mathrm{MSE}&=\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]\\ \mathrm{MV}&=\operatorname*{\mathbb{E}}_{D_{r},\theta}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{SDV}&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\theta|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]\end{split}\quad\quad\begin{split}\mathrm{RDV}&=\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x;D_{s})]\\ \mathrm{SDB}&=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|D_{r}}[f_{\theta}(x)]]\\ \mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]].\end{split}$ (4)
$f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor for real data, $\theta\sim p(\theta|D_{r})$ is a single sample from the distribution of the generator parameters, $D_{s}\sim p(D_{s}|\theta)$ is a single sample of the synthetic data generating process, and $f_{\theta}$ is the optimal predictor for the synthetic data generating process with parameters $\theta$. All random quantities are implicitly conditioned on $x$. The proofs of all theorems are in Appendix A.
@@
    Increasing the number of synthetic datasets $m$ reduces the impact of MV and SDV. This means that high-variance models, like interpolating decision trees or 1-nearest neighbours, benefit the most from multiple synthetic datasets.
@@
    If $\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]$ is invariant to the size of the synthetic dataset, the only term that can change with the size of the synthetic dataset is MV. This means that the possible benefit of generating a single large synthetic dataset is smaller than generating many synthetic datasets, which can also reduce the impact of SDV.
  4. 4.
    The quality metrics for a synthetic data generator should include metrics that compare the distribution of the synthetic data to the real distribution, instead of just comparing a synthetic dataset to the real dataset. An example of the former is the downstream performance of predictive models, and examples of the latter are comparisons between the marginals of the real and synthetic datasets.
@@
    $\begin{split}\mathrm{MV}&=\operatorname*{Var}_{D_{r}}[g(x;D_{r})]\\ \mathrm{SDV}&=\mathrm{RDV}=\mathrm{SDB}=0.\end{split}\quad\quad\begin{split}\mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-g(x;D_{r})]\\ \phantom{x}\end{split}$ (5)
@@
    Return the real data: $D_{s}|D_{r}=D_{r}$ deterministically.333 Theorem 2.2 applies in this scenario even with multiple synthetic datasets, as the deterministically identical synthetic datasets are independent as random variables.
    $\begin{split}\mathrm{RDV}&=\operatorname*{Var}_{D_{r}}[g(x;D_{r})]\\ \mathrm{MV}&=\mathrm{SDV}=\mathrm{SDB}=0.\end{split}\quad\quad\begin{split}\mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-g(x;D_{r})]\\ \phantom{x}\end{split}$ (6)
@@ 2.3 Estimating the effect of Multiple Synthetic Datasets
From (8), we see that $\mathrm{MV}+\mathrm{SDV}$ is the maximum decrease in MSE that can be obtained from multiple synthetic datasets. This means that $2(\mathrm{MSE}_{1}-\mathrm{MSE}_{2})$, or the linear regression estimate from (10), can be used as a diagnostic to check whether generating multiple synthetic datasets is worthwhile.
@@ 2.4 Differentially Private Synthetic Data Generators
The second possibility is generating all synthetic datasets based on a single application of a DP mechanism. Specifically, a noisy summary $\tilde{s}$ of the real data is released under DP. The parameters $\theta_{1:m}$ are then sampled i.i.d. conditional on $\tilde{s}$, and the synthetic datasets are sampled conditionally on the $\theta_{1:m}$. This setting includes algorithms that release a posterior distribution under DP, and use the posterior to generate synthetic data, like the NAPSU-MQ algorithm (Räisä et al., 2023b) and DP variational inference (DPVI) (Jälkö et al., 2017, 2021).444 In DPVI, $\tilde{s}$ would be the variational approximation to the posterior.
@@
$\begin{split}\mathrm{MSE}&=\frac{1}{m}\mathrm{MV}+\frac{1}{m}\mathrm{SDV}+\mathrm{RDV}+\mathrm{DPVAR}+\,(\mathrm{SDB}+\mathrm{MB})^{2}+\operatorname*{Var}_{y}[y],\end{split}$ (11)
@@
$\begin{split}\mathrm{MSE}&=\operatorname*{\mathbb{E}}_{y,D_{r},\tilde{s},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]\\ \mathrm{MV}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{SDV}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}\operatorname*{Var}_{\theta|\tilde{s}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{RDV}&=\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x;D_{s})].\end{split}\quad\quad\begin{split}\mathrm{DPVAR}&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\tilde{s}}[g(x;D_{s})]\\ \mathrm{SDB}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|\tilde{s}}[f_{\theta}(x)]]\\ \mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]]\\ \phantom{x}\end{split}$ (12)

### v3

@@ Abstract
Recent studies have highlighted the benefits of generating multiple synthetic datasets for supervised learning, from increased accuracy to more effective model selection and uncertainty estimation. These benefits have clear empirical support, but the theoretical understanding of them is currently very light. We seek to increase the theoretical understanding by deriving bias-variance decompositions for several settings of using multiple synthetic datasets, including differentially private synthetic data. Our theory yields a simple rule of thumb to select the appropriate number of synthetic datasets in the case of mean-squared error and Brier score. We investigate how our theory works in practice with several real datasets, downstream predictors and error metrics. As our theory predicts, multiple synthetic datasets often improve accuracy, while a single large synthetic dataset gives at best minimal improvement, showing that our insights are practically relevant.
@@ 1 INTRODUCTION
In supervised learning, van Breugel et al. (2023a) proposed an ensemble method of generative models. They propose generating multiple synthetic datasets independently from an arbitrary generative model, training a predictive model separately on each synthetic dataset, and ensembling these models by averaging their predictions. They call this a deep generative ensemble (DGE). We drop the word “deep” in this paper and use the term generative ensemble, as we do not require the generator to be deep in any sense. The DGE was empirically demonstrated to be beneficial in several ways by van Breugel et al. (2023a), including predictive accuracy, model evaluation, model selection, and uncertainty estimation. Followup work has applied DGEs to improve model evaluation under distribution shifts and for small subgroups (van Breugel et al., 2023b).
@@
$\underbrace{\operatorname*{\mathbb{E}}_{D,y}[(y-g)^{2}]}_{\mathrm{MSE}}=\underbrace{(f(x)-\operatorname*{\mathbb{E}}_{D}[g])^{2}}_{\mathrm{Bias}}+\underbrace{\operatorname*{Var}_{D}[g]}_{\mathrm{Variance}}+\underbrace{\operatorname*{Var}_{y}[y]}_{\mathrm{Noise}},$ (1)
where we have shortened $g=g(x;D)$ and $f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor. $x$ is considered fixed, so all the random quantities in the decomposition are implicitly conditioned on $x$. While MSE is typically only used with regression problems, the decomposition also applies to the Brier score (Brier, 1950) in classification, which is simply the MSE of class probabilities.
@@
    We derive a bias-variance decomposition for the MSE or Brier score of generative ensembles under an i.i.d. assumption in Theorem 2.2. This decomposition is simply a sum of interpretable terms, which makes it possible to predict how various choices affect the error.
@@
    We derive several practical considerations from our decomposition, including a simple rule of thumb to select the number of synthetic datasets in Section 2.3. In summary, 2 synthetic datasets give 50% of the potential benefit from multiple synthetic datasets, 10 give 90% of the benefit and 100 give 99% of the benefit. This also applies to bagging ensembles like random forests, which is likely to be of independent interest. The benefit is a result of reduced variance, so the theory predicts high-variance models to receive the largest benefit.
@@
    We evaluate the performance of a generative ensemble on several datasets, downstream prediction algorithms, and error metrics in Section 3. The results show that out theory applies in practice: multiple synthetic datasets generally decrease all of the error metrics, and our rule of thumb makes accurate predictions when the error can be accurately estimated. In contrast, a single large synthetic dataset provides small benefits at best, and can even increase error.
@@ 1.1 Related Work
Ueda & Nakano (1996) were the first to study the MSE bias–variance decomposition with ensembles, and Gupta et al. (2022); Wood et al. (2023) later extended the ensemble decomposition to other losses. All of these also apply to generative ensembles, but they only provide limited insight for them, as they do not separate the synthetic data generation-related terms from the downstream-related terms.
@@ 2 BIAS-VARIANCE DECOMPOSITIONS FOR GENERATIVE ENSEMBLES
In this section, we make our main theoretical contributions. We start by defining our setting in Section 2.1. We then derive a bias-variance decomposition for generative ensembles in Section 2.2, and a simple rule of thumb that can be used to select the number of synthetic datasets in Section 2.3. The first decomposition does not apply to some differentially private synthetic data generation methods, so we generalise the decomposition to apply to those in Section 2.4. We also present decompositions that apply non-i.i.d.  settings and to Bregman divergences in Appendices B and C, but they are not as informative as the others.
@@ 2.1 Problem Setting
We assume that each synthetic dataset $D_{s}^{i}$ is generated from a generator with parameters $\theta_{i}$, and that each generator is run with an independent random seed, so generations are independent given the parameters $\theta_{1:m}$. We also assume that each generator is trained on the real data $D_{r}$ with an independent random seed, and there are no other dependencies between the training processes besides the real data, so the $\theta_{i}$ are i.i.d. given $D_{r}$. This is how synthetic datasets are sampled in DGE (van Breugel et al., 2023a), where $\theta_{i}$ are the parameters of the generative neural network from independent training runs. This also encompasses bootstrapping, where $\theta_{i}$ would be the real dataset [^2] The $\theta_{i}$ random variables are i.i.d. if they are deterministically equal., and $p(D_{s}|\theta_{i})$ is the bootstrapping.
We will also consider a more general setting that applies to some differentially private synthetic data generators that do not fit into this setting in Section 2.4. In Appendix B, we consider settings without the i.i.d. assumptions.
@@ 2.2 Mean-squared Error Decomposition
Next, we present our main theorem. With the conditional i.i.d. assumptions detailed in Section 2.1, all of the $\theta_{i}|D_{r}$ distributions are identical and independent, so we use $\theta$ as a shorthand for a random variable with this distribution in this section. Similarly, the $D_{s}^{i}|\theta_{i}$ distribution only depends on $\theta_{i}$, but not $i$ with these assumptions, so we use $D_{s}|\theta$ as a shorthand.
@@
$\begin{split}\mathrm{MSE}&=\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]\\ \mathrm{MV}&=\operatorname*{\mathbb{E}}_{D_{r},\theta}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{SDV}&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\theta|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{RDV}&=\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x;D_{s})]\\ \mathrm{SDB}&=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|D_{r}}[f_{\theta}(x)]]\\ \mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]].\end{split}$ (4)
$f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor for real data, $\theta\sim p(\theta|D_{r})$ is a single sample from the identical distribution of the generator parameters $\theta_{i}$, $D_{s}\sim p(D_{s}|\theta)$ is a single sample of the synthetic data generating process given $\theta$, and $f_{\theta}$ is the optimal predictor for the synthetic data generating process with parameters $\theta$. All random quantities are implicitly conditioned on $x$. The proofs of all theorems are in Appendix A.
To intuitively explain what each term measures, we split the whole generative ensemble pipeline into three steps:
  1. 1.
    Sample real data $D_{r}$,
  2. 2.
    given $D_{r}$, train generative model, resulting in $\theta$,
  3. 3.
    given $\theta$, sample synthetic data $D_{s}$, train downstream model $g$ and make a prediction on $x$.
In the full generative ensemble, steps 2 and 3 are repeated $m$ times.
MV measures the variance of step 3, averaged over the randomness in steps 1 and 2. SDV measures the variance of the average prediction from step 3 over the randomness in step 2, and also averages over the randomness in step 1. RDV measures the variance of the average prediction from the combination of steps 2 and 3, over the randomness in step 1.
$f_{\theta}$ represents the optimal predictor if the synthetic data generation in step 3 were the real data generating process. SDB measures how much the average $f(\theta)$, over the randomness in step 2, differs from the actual optimal predictor $f$ on the real data, and averages the difference over the randomness in step 1. MB measures how much the downstream model’s average predictions, over the randomness in step 3, differ from $f_{\theta}$, and averages the difference over the randomness in steps 1 and 2. $\operatorname*{Var}_{y}[y]$ is the noise term also present in the classical bias-variance decomposition in (1), which represents inherent randomness in the prediction problem that cannot be removed by any predictor.
Note that the MV, SDV, RDV, SDB, MB and $\operatorname*{Var}_{y}[y]$ terms do not depend on $m$. Changing $m$ only affects the impact of MV and SDV on MSE, as seen in (3).
Multidimensional $y$. For a multidimensional $y\in\mathbb{R}^{d}$, we have
$\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}\big{[}||y-\hat{g}||_{2}^{2}\big{]}=\sum_{j=1}^{d}\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}\big{[}(y_{j}-\hat{g_{j}})^{2}\big{]},$ (5)
where we have shortened $\hat{g}=\hat{g}(x;D_{s}^{1:m})$ and $y_{j}$ and $\hat{g}_{j}$ index over the dimensions. Since the right hand side is a sum of one-dimensional MSEs, our theory can also be applied to multi-dimensional $y$.
Relation to Brier Score. The Brier score (Brier, 1950) can be defined in two ways for binary classification. The first way to is pick a class, here class 0, and define
$\mathrm{BS}_{1}=\operatorname*{\mathbb{E}}_{y,D}[(y_{0}-g_{0}(x;D))^{2}],$ (6)
where $y_{0}$ is the indicator for the true class being 0, and $g_{0}(x;D)$ is the predicted probability of class 0. The second way is to define
$\mathrm{BS}_{2}=\operatorname*{\mathbb{E}}_{y,D}[||y-g(x;D)||_{2}^{2}],$ (7)
where $y$ is a one-hot-encoding of the true class, and $g(x;D)$ is the vector of predicted probabilities. The second definition also extends to problems with more than two classes. For binary classification, $\mathrm{BS}_{2}=2\cdot\mathrm{BS}_{1}$ since probabilities sum to 1, so it does not matter which class is picked for the first definition.
The first definition is an MSE, and the second is a sum of one-dimensional MSEs, so our theory applies to both, including multiclass problems through the extension to multidimensional $y$.
Practical Considerations
@@
    Increasing the number of synthetic datasets $m$ reduces the impact of MV and SDV. This means that high-variance models, like interpolating decision trees or 1-nearest neighbours, benefit the most from multiple synthetic datasets. The size of this benefit can be empirically estimated using the prediction rule of Section 2.3.
@@
    The quality metrics for a synthetic data generator should include metrics that compare the distribution of the synthetic [^3] Specifically, the conditional distribution of synthetic data given real data. data to the real distribution, instead of just comparing a synthetic dataset to the real dataset. Examples of the latter are all metrics that return the optimal value for a generator that just returns the real data, like comparisons between the marginals of the real and synthetic datasets, comparisons of downstream ML performance, and metrics like density/coverage (Naeem et al., 2020). One example of the former is the authenticity metric of (Alaa et al., 2022), though it only attempts to capture how well the generator generalises instead of copying the real data.
@@
    $\begin{split}\mathrm{MV}&=\operatorname*{Var}_{D_{r}}[g(x;D_{r})],\,\,\,\,\mathrm{MB}=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-g(x;D_{r})]\\ \mathrm{SDV}&=\mathrm{RDV}=\mathrm{SDB}=0.\end{split}$
@@
    Return the real data: $D_{s}|D_{r}=D_{r}$ deterministically.444 Theorem 2.2 applies in this scenario even with multiple synthetic datasets, as the deterministically identical synthetic datasets are independent as random variables.
    $\begin{split}\mathrm{RDV}&=\operatorname*{Var}_{D_{r}}[g(x;D_{r})],\,\,\,\mathrm{MB}=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-g(x;D_{r})]\\ \mathrm{MV}&=\mathrm{SDV}=\mathrm{SDB}=0.\end{split}$
@@ 2.3 Estimating the effect of Multiple Synthetic Datasets
From (9), we see that $\mathrm{MV}+\mathrm{SDV}$ is the maximum reduction in MSE that can be obtained from multiple synthetic datasets. This means that $2(\mathrm{MSE}_{1}-\mathrm{MSE}_{2})$, or the linear regression estimate from (11), can be used as a diagnostic to check whether generating multiple synthetic datasets is worthwhile.
@@ 2.4 Differentially Private Synthetic Data Generators
The second possibility is generating all synthetic datasets based on a single application of a DP mechanism. Specifically, a noisy summary $\tilde{s}$ of the real data is released under DP. The parameters $\theta_{1:m}$ are then sampled i.i.d. conditional on $\tilde{s}$, and the synthetic datasets are sampled conditionally on the $\theta_{1:m}$. This setting includes algorithms that release a posterior distribution under DP, and use the posterior to generate synthetic data, like the NAPSU-MQ algorithm (Räisä et al., 2023b) and DP variational inference (DPVI) (Jälkö et al., 2017, 2021).555 In DPVI, $\tilde{s}$ would be the variational approximation to the posterior.
@@
$\begin{split}\mathrm{MSE}&=\frac{1}{m}\mathrm{MV}+\frac{1}{m}\mathrm{SDV}+\mathrm{RDV}+\mathrm{DPVAR}\\ &+(\mathrm{SDB}+\mathrm{MB})^{2}+\operatorname*{Var}_{y}[y],\end{split}$ (12)
@@
$\begin{split}\mathrm{MSE}&=\operatorname*{\mathbb{E}}_{y,D_{r},\tilde{s},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]\\ \mathrm{MV}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{SDV}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}\operatorname*{Var}_{\theta|\tilde{s}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{RDV}&=\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x;D_{s})]\end{split}$ (13)
$\begin{split}\mathrm{DPVAR}&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\tilde{s}}[g(x;D_{s})]\\ \mathrm{SDB}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|\tilde{s}}[f_{\theta}(x)]]\\ \mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]].\end{split}$ (14)